# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 6.3 Recursion
In this section, we show what information can be extracted from the recursive relationship between the sum of top eigenvalues, that is Lemmas 6.3 and 6.5. The proof is just a simple analysis on sequences.
@@
Let $n,k$ be integers, $0<s<q$ real numbers, and let $\lambda_{1}\geq\dots\geq\lambda_{k}\geq 0$. For every real number $T$, define $S_{T}=\sum_{\lambda_{i}\geq T}\lambda_{i}$. Assume that $\lambda_{1}+\dots+\lambda_{k}\leq n^{1+s}$, and for every $T\geq n^{1-q}$,
$S_{T}^{2}\leq 4nS_{\frac{T^{2}}{2n}}.$
Then for every $H\leq n$,
$\sum_{\lambda_{i}\leq H}\lambda_{i}^{2}\leq\frac{16q}{q-s}n^{1+s/q}H^{1-s/q}.$
@@
Let $T_{0}=n^{1-2q}/2$, and recursively define $T_{i+1}=(2nT_{i-1})^{1/2}$ for $i=0,1,\dots$. Observe that
$\frac{2n}{T_{i+1}}=\left(\frac{2n}{T_{i}}\right)^{1/2}.$
The sequence $T_{0},T_{1},\dots$ is monotone increasing, and we have $T_{1}=(2nT_{0})^{1/2}\geq n^{1-q}$, so we get that
$S_{T_{i+1}}^{2}\leq 4nS_{T_{i}}$
for all $i\geq 0$. We prove by induction on $i$ that
$S_{T_{i}}\leq 4n\left(\frac{2n}{T_{i+1}}\right)^{s/q}.$
For the base case $i=0$, this is true as
$S_{T_{0}}\leq\sum_{i=1}^{k}\lambda_{i}\leq n^{1+s}=n\left(\frac{n}{n^{1-q}}\right)^{s/q}\leq 4n\left(\frac{2n}{T_{1}}\right)^{s/q}.$
Now let $i>0$, then
$S_{T_{i+1}}\leq\sqrt{4nS_{T_{i}}}\leq\sqrt{16n^{2}\left(\frac{2n}{T_{i+1}}\right)^{s/q}}=4n\left(\frac{2n}{T_{i+2}}\right)^{s/q}.$
It is easy to see that $\lim_{i\rightarrow\infty}T_{i}=2n$, so for every $T\in[T_{0},n]$, there exists an index $i$ such that $T_{i}\leq T<T_{i+1}$. As $S_{T}$ is decreasing in $T$, we have
$S_{T}\leq S_{T_{i}}\leq 4n\left(\frac{2n}{T_{i+1}}\right)^{s/q}\leq 4n\left(\frac{2n}{T}\right)^{s/q}.$
This inequality also holds for $T\leq T_{0}=n^{1-2p}/2$ as
$S_{T}\leq S_{T_{0}}\leq n^{1+s}\leq 4n\left(\frac{2n}{T}\right)^{s/q}.$
Hence, we can write
$\sum_{\lambda_{i}\leq H}\lambda_{i}^{2}\leq\int_{0}^{H}2S_{t}dt\leq\int_{0}^{H}8n\left(\frac{2n}{t}\right)^{s/q}dt=\frac{8\cdot 2^{s/q}}{1-s/q}n^{1+s/p}H^{1-s/q}.$
In the last equality, we use that $s/q<1$. Moreover, the first inequality follows as
$\sum_{\lambda_{i}\leq H}\lambda_{i}^{2}=\sum_{\lambda_{i}}\int_{0}^{\lambda_{i}}2tdt=\int_{0}^{H}2N_{t}tdt\leq\int_{0}^{H}2S_{t}dt,$
where $N_{t}=\#\{i:\lambda_{i}\geq t\}$. ∎
7 Densification — Phase 0
In this section, we prove that every graph of average degree $d$ and smallest eigenvalue $\lambda_{n}$ contains a subgraph on $d$ vertices of density $\Omega(1/|\lambda_{n}|)$.
@@ 7 Densification — Phase 0
Let $G$ be an $n$-vertex graph of average degree $d$ with smallest eigenvalue $\lambda_{n}$ satisfying $|\lambda_{n}|\leq\frac{1}{2}d^{1/2}$. Then $G$ contains an induced subgraph on $d$ vertices with edge density $\Omega(1/|\lambda_{n}|)$.
@@
Let $x$ be a vertex with at least $d$ neighbours, and let $S$ be a set of exactly $d$ neighbours of $x$. Define the vector $v\in\mathbb{R}^{n}$ such that $v(x)=1$, $v(y)=\frac{\lambda_{n}}{d}$ if $y\in Y$, and $v(z)=0$ if $z\not\in S\cup\{x\}$. Then
$v^{T}Av=2\lambda_{n}+2\frac{\lambda_{n}^{2}}{d^{2}}e(G[S]).$
On the other hand
$v^{T}v=1+\frac{\lambda_{n}^{2}}{d}\leq\frac{3}{2}.$
Therefore, we obtain that
$2\lambda_{n}+2\frac{\lambda_{n}^{2}}{d^{2}}e(G[S])=v^{T}Av\geq\lambda_{n}v^{T}v\geq\frac{3}{2}\lambda_{n}.$
From this,
$e(G[S])\geq\frac{d^{2}}{4|\lambda_{n}|},$
showing that $G[S]$ is the desired subgraph. ∎
8 Densification — Phase 1
Lemma 6.7, combined with earlier results, has a number of powerful consequences. First, we use it to show that a graph with large smallest eigenvalue (or small surplus) contains a subgraph of positive density. We prove this via the following density increment argument. We apply Lemma 6.7 to show that the cubic sum of eigenvalues is large, which in turn coincides with six times the number of triangles. But if a graph has too many triangles, it means that there is a vertex, whose neighborhood is dense, so we pass to this neighborhood, and repeat. The main step of this argument is presented in the next lemma.
Lemma 8.1.
Let $s\in(0,1/6)$, $C>2$, and let $G$ be a $n$-vertex graph with edge density $p>n^{-1/2}$, $\Delta(G)\leq Cpn$, and smallest eigenvalue $\lambda_{n}$ satisfying $|\lambda_{n}|\leq n^{s}$. Then $G$ has a subgraph on at least $pn$ vertices of edge density at least $c_{0}p^{6s}$ for some $c_{0}=c_{0}(s,C)>0$.
Proof.
Let $m=p\binom{n}{2}$ denote the number of edges. Let $\lambda_{1}\geq\dots\geq\lambda_{n}$ be the eigenvalues of $G$, then
$\sum_{\lambda_{i}>0}\lambda_{i}=\sum_{0<\lambda_{i}}|\lambda_{i}|\leq n|\lambda_{n}|\leq n^{1+s}.$
@@ 8 Densification — Phase 1
for every $T\geq 2n^{1/2+s}$. So applying Lemma 6.7 with $q=1/3>s$ and the sequence of positive eigenvalues, we get that for every $H\leq n$,
$\sum_{0<\lambda_{i}\leq H}\lambda_{i}^{2}\leq cn^{1+3s}H^{1-3s}$
for some $c>1$ depending only on $s$. Write $u:=3s<1/2$, and set $H:=(8c)^{-1/(1-u)}p^{1/(1-u)}n$, then $H\leq n$ and the right-hand-side equals $pn^{2}/8$. On the other hand, we have
$\sum_{\lambda_{i}<0}\lambda_{i}^{2}\leq n\lambda_{n}^{2}\leq n^{1+2s}<pn^{2}/8,$
@@
$\sum_{\lambda_{i}\leq H}\lambda_{i}^{2}\leq pn^{2}/4.$
[…]
$\sum_{\lambda_{i}>H}\lambda_{i}^{2}\geq 2m-\sum_{\lambda_{i}\leq H}\lambda_{i}^{2}\geq pn^{2}/2.$
@@
$6N=\sum_{i=1}^{n}\lambda_{i}^{3}\geq H\sum_{\lambda_{i}\geq H}\lambda_{i}^{2}-\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\geq Hpn^{2}/2-n|\lambda_{n}|^{3}\geq c^{\prime}p^{\frac{2-u}{1-u}}n^{3}-n^{1+3s}$
for some constant $c^{\prime}>0$ depending only on $s$. Here, using that $s<1/6$, $u<1/2$ and $p>n^{-1/2}$, we have $n^{1+3s}<\frac{c^{\prime}}{2}p^{\frac{2-u}{1-u}}n^{3}$ for sufficiently large $n$. Hence, we get
@@
Counting triangles by vertices, we observe that there is a vertex $v\in V(G)$ whose neighborhood $X$ contains at least $\frac{3N}{n}$ edges. Here, note that $|X|\leq\Delta(G)\leq Cpn$. Let $X^{\prime}$ be an arbitrary $\min\{Cpn,n\}$ element subset of $V(G)$ containing $X$, then the edge density of $G[X^{\prime}]$ is
$\frac{3N/n}{\binom{|X^{\prime}|}{2}}\geq\frac{c^{\prime}p^{\frac{2-u}{1-u}}n^{2}}{2C^{2}p^{2}n^{2}}=\frac{c^{\prime}}{2C^{2}}p^{\frac{u}{1-u}}.$
As u/(1u)=3s/(13s)<6su/(1-u)=3s/(1-3s)<6s, this finishes the proof. ∎
@@
Let $C>2$, let $G$ be an $n$-vertex graph of average degree $d$ such that $\operatorname{surp}(G)\leq 0.01dn$. Then either
@@
  • (ii)
    $G$ contains a subgraph on at least $n/C$ vertices of average degree at least $0.2Cd$.
@@
Let $X\subset V(G)$ be the set of vertices of degree more than $Cd$, then $|X|\leq n/C$. Let $Y=V(G)\setminus X$. The maximum degree of $G[Y]$ is at most $Cd$, so if $G[Y]$ has average degree at least $d/4$, then (i) holds. Otherwise, there are at most $nd/8$ edges in $G[X]$. As $\operatorname{surp}(G)\leq 0.1dn$, we have $e(G[X,Y])\leq\frac{dn}{4}+\operatorname{surp}(G)\leq 0.26dn$. Hence, we must have that $G[Y]$ has at least dn/2dn/80.26dn>0.1dndn/2-dn/8-0.26dn>0.1dn edges. Let $X_{0}$ be an arbitrary $n/C$ sized subset of $V(G)$ containing $X$, then $X_{0}$ has average degree at least $0.2Cd$, so $G[X_{0}]$ satisfies the requirements. ∎
@@
Let $s\in(0,1/6)$, $\varepsilon\in(0,1/2)$, then there exists $\rho=\rho(s,\varepsilon)>0$ and $c_{1}=c_{1}(s,\varepsilon)>0$ such that the following holds for every $n>n_{0}(s,\varepsilon)$. Let $G$ be an $n$-vertex graph with edge density $p$ and smallest eigenvalue $\lambda_{n}$ such that $p>n^{-\rho}$ and $|\lambda_{n}|\leq n^{s}$. Then $G$ has a subgraph on at least $n^{1-\varepsilon}$ vertices with edge density at least $c_{1}$.
We may choose $\rho=\alpha\varepsilon$ with some decreasing function $\alpha=\alpha(s)>0$.

### v2

@@ 6.3 Recursion
In this section, we show how to use the inequality $S_{T}^{2}\leq 4nS_{\frac{T^{2}}{2n}}$ to gain insight into the global structure of the spectrum. This section can be summed up in the motto “large eigenvalues carry all the mass in the second moment”.
@@
Let $n$ be an integer, $0<\gamma<q<1$ be fixed parameters, and let $\lambda_{1}\geq\dots\geq\lambda_{n}$ be the eigenvalues of an $n$-vertex graph $G$. Assume that $\sum_{\lambda_{i}>0}\lambda_{i}\leq n^{1+\gamma}$ and $S_{T}^{2}\leq 4nS_{\frac{T^{2}}{2n}}$, for every $T\geq 2n^{1-q}$. Then for every $\kappa\in[0,1]$,
$\sum_{0\leq\lambda_{i}\leq\kappa n}\lambda_{i}^{2}\leq 50\kappa^{1-\gamma/q}n^{2}.$ (6)
@@
First, observe that if $\kappa\leq 2n^{-q}$, the statement is easy to prove. Indeed, using that $0<\gamma<q<1$,
$\sum_{0\leq\lambda_{i}\leq\kappa n}\lambda_{i}^{2}\leq\kappa n\sum_{\lambda_{i}>0}\lambda_{i}\leq\kappa n^{2+\gamma}\leq\kappa n^{2}\cdot(2/\kappa)^{\gamma/q}\leq 2^{\gamma/q}\kappa^{1-\gamma/q}n^{2}\leq 50\kappa^{1-\gamma/q}n^{2}.$
In the rest of the proof, we assume $\kappa>2n^{-q}$. Define an increasing sequence $\kappa_{-1},\kappa_{0},\kappa_{1},\dots$ by $\kappa_{-1}=2n^{-2q}$ and $\kappa_{i}=\sqrt{2\kappa_{i-1}}$ for $i\geq 0$. Then $\kappa_{i}=2n^{-q/2^{i}}$ for $i\geq-1$, explaining our reason to start the indexing from 1-1. We also highlight that $\kappa>\kappa_{-1}$.
First, we show inductively that $S_{\kappa_{i}n}\leq 8\kappa_{i}^{-\gamma/2q}n$. For i=1i=-1 this is straightforward, since
$S_{\kappa_{-1}n}\leq\sum_{\lambda_{j}>0}\lambda_{j}\leq n^{1+\gamma}\leq 8\kappa_{-1}^{-\gamma/2q}n.$
For $i\geq 0$, we have
$S_{\kappa_{i}n}\leq\sqrt{4nS_{\kappa_{i}^{2}n/2}}=\sqrt{4nS_{\kappa_{i-1}n}}\leq\sqrt{4n\cdot 8\kappa_{i-1}^{-\gamma/2q}n}=\sqrt{32(\kappa_{i}^{2}/2)^{-\gamma/2q}n^{2}}\leq 8\kappa_{i}^{-\gamma/2q}n,$
where in the first inequality we use that $\kappa_{i}n\geq 2n^{1-q}$ for $i\geq 0$, so $S_{T}^{2}\leq 4nS_{\frac{T^{2}}{2n}}$ holds for $T=\kappa_{i}n$, and in the second inequality we use the induction hypothesis for $S_{\kappa_{i-1}n}$.
Next, we prove (6). We may assume $\kappa^{1-\gamma/q}\leq 1/2$, since the statement is otherwise trivial. Let $p$ be the largest integer such that $\kappa_{p}\leq\kappa$, for which we have $\kappa<\sqrt{2\kappa_{p}}$, i.e. $\kappa_{p}>\kappa^{2}/2$. Then, we can decompose the sum $\sum_{0\leq\lambda_{i}\leq\kappa n}\lambda_{i}^{2}$ as follows
$\sum_{0\leq\lambda_{i}\leq\kappa n}\lambda_{i}^{2}$ $=\underbrace{\sum_{\kappa_{p}n<\lambda_{i}\leq\kappa n}\lambda_{i}^{2}}_{\Sigma_{p}}+\underbrace{\sum_{\kappa_{p-1}n<\lambda_{i}\leq\kappa_{p}n}\lambda_{i}^{2}}_{\Sigma_{p-1}}+\cdots+\underbrace{\sum_{0<\lambda_{i}\leq\kappa_{-1}n}\lambda_{i}^{2}}_{\Sigma_{-2}}.$
We can bound $\Sigma_{p}$ as
$\Sigma_{p}\leq\kappa n\cdot S_{\kappa_{p}n}\leq\kappa n\cdot 8\kappa_{p}^{-\gamma/2q}n\leq 16\kappa^{1-\gamma/q}n^{2},$
where in the last step we used that $\kappa_{p}>\kappa^{2}/2$. Similarly, for any $1\leq\ell\leq p-1$, we have
$\Sigma_{\ell-1}\leq\kappa_{\ell}n\cdot S_{\kappa_{\ell-1}n}\leq\kappa_{\ell}n\cdot 8\kappa_{\ell-1}^{-\gamma/2q}n\leq 16\kappa_{\ell}^{1-\gamma/q}n^{2},$
where we again used that $\kappa_{\ell-1}=\kappa_{\ell}^{2}/2$. Finally, using that $\kappa_{-1}\leq 2n^{-q}$, we have
$\Sigma_{-2}\leq\kappa_{-1}n\sum_{\lambda_{i}>0}\lambda_{i}\leq\kappa_{-1}n^{2+\gamma}\leq 2\kappa_{-1}^{1-\gamma/q}n^{2}.$
Combining all of this, we obtain
$\sum_{0\leq\lambda_{i}\leq\kappa n}\lambda_{i}^{2}\leq 16\kappa^{1-\gamma/q}n^{2}+\sum_{\ell\leq p}16\kappa_{\ell}^{1-\gamma/q}n^{2}+2\kappa_{-1}^{1-\gamma/q}n^{2}.$
We note that $\kappa_{\ell}^{1-\gamma/q}\leq 2^{-(p-\ell)}\kappa^{1-\gamma/q}$, which follows by simple induction and our assumption that $\kappa^{1-\gamma/q}<1/2$. Thus, the right-hand-side of the above inequality is less than $50\kappa^{1-\gamma/q}n$. ∎
7 Densification — Phase 1
Lemma˜6.9, combined with earlier results, has a number of powerful consequences. One of them asserts that a graph with large smallest eigenvalue (or small surplus) contains a large subgraph of positive density. We prove this via the following density increment argument. We apply Lemma˜6.9 to show that the cubic sum of eigenvalues is large, which in turn coincides with six times the number of triangles. But if a graph has too many triangles, then some vertex has a very dense neighbourhood. So we can repeatedly pass to this neighbourhood to eventually obtain a large subgraph with constant density. The main step of this argument is presented in the next lemma.
@@ 7 Densification — Phase 1
Let $\gamma\in(0,1/6)$, $C>2$, and let $G$ be a $n$-vertex graph with edge density $p>n^{-1/2}$, $\Delta(G)\leq Cpn$, and smallest eigenvalue $\lambda_{n}$ satisfying $|\lambda_{n}|\leq n^{\gamma}$. Then $G$ has a subgraph on at least $pn$ vertices of edge density at least $c_{0}p^{2\gamma/(1-4\gamma)}$ for some $c_{0}=c_{0}(\gamma,C)>0$.
@@
We may and will assume $n$ is sufficiently large. Let $m=p\binom{n}{2}$ denote the number of edges. Let $\lambda_{1}\geq\dots\geq\lambda_{n}$ be the eigenvalues of $G$, then
$\sum_{\lambda_{i}>0}\lambda_{i}=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq n|\lambda_{n}|\leq n^{1+\gamma}.$
@@
for every $T\geq 2n^{1/2+\gamma}$. So applying Lemma 6.9 with $q=1/2-\gamma>\gamma$ and the sequence of positive eigenvalues, we get that for every $\kappa\in[0,1]$,
$\sum_{0<\lambda_{i}\leq\kappa n}\lambda_{i}^{2}\leq c\kappa^{1-\gamma/q}n^{2}$
for an absolute constant $c>1$. Write $u:=\gamma/q<1/2$ and set $\kappa:=(8c)^{-1/(1-u)}p^{1/(1-u)}$. Then $\kappa\leq 1$ and the right-hand-side equals $pn^{2}/8$. On the other hand, we have
$\sum_{\lambda_{i}<0}\lambda_{i}^{2}\leq n\lambda_{n}^{2}\leq n^{1+2\gamma}<pn^{2}/8,$
@@
$\sum_{\lambda_{i}\leq\kappa n}\lambda_{i}^{2}\leq pn^{2}/4.$
But $\sum_{i=1}^{n}\lambda_{i}^{2}=\|A\|_{F}^{2}=2m$, so we conclude that
$\sum_{\lambda_{i}>\kappa n}\lambda_{i}^{2}\geq 2m-\sum_{\lambda_{i}\leq H}\lambda_{i}^{2}\geq pn^{2}/2.$
@@
$6N=\sum_{i=1}^{n}\lambda_{i}^{3}\geq\kappa n\sum_{\lambda_{i}>\kappa n}\lambda_{i}^{2}-\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\geq\kappa pn^{3}/2-n|\lambda_{n}|^{3}\geq c^{\prime}p^{\frac{2-u}{1-u}}n^{3}-n^{1+3\gamma}$
for some constant $c^{\prime}>0$ depending only on $\gamma$. Here, using that $\gamma<1/6$, $u<1/2$ and $p>n^{-1/2}$, we have $n^{1+3\gamma}<\frac{c^{\prime}}{2}p^{\frac{2-u}{1-u}}n^{3}$. Hence, we get
@@
Counting triangles by vertices, we observe that there is a vertex $v\in V(G)$ whose neighbourhood $X$ contains at least $\frac{3N}{n}$ edges. In addition, $|X|\leq\Delta(G)\leq Cpn$. Now, let $X^{\prime}$ be an arbitrary superset of $X$ containing $\max(pn,|X|)$ elements. Then, the edge density of $G[X^{\prime}]$ is
$\frac{3N/n}{\binom{|X^{\prime}|}{2}}>\frac{6N/n}{|X^{\prime}|^{2}}\geq\frac{c^{\prime}p^{\frac{2-u}{1-u}}n^{2}/2}{C^{2}p^{2}n^{2}}=\frac{c^{\prime}}{2C^{2}}p^{\frac{u}{1-u}}.$
As $\frac{u}{1-u}=\frac{\gamma}{q-\gamma}=\frac{2\gamma}{1-4\gamma}$, this finishes the proof. ∎
@@
Let $C>2$ and $G$ be an $n$-vertex graph of average degree $d$ with $\operatorname{surp}(G)\leq\frac{dn}{100}$. Then either
@@
    $G$ contains a subgraph on $n/C$ vertices of average degree at least $Cd/5$.
  • (ii)
@@
Let $X\subset V(G)$ be the set of vertices of degree more than $Cd$, then $|X|\leq n/C$. Let $Y=V(G)\setminus X$. The maximum degree of $G[Y]$ is at most $Cd$, so if $G[Y]$ has average degree at least $d/4$, then (ii) holds. Otherwise, $G[Y]$ has at most $nd/8$ edges. Moreover, $e(G[X,Y])\leq\frac{e(G)}{2}+\operatorname{surp}(G)\leq\frac{dn}{4}+\frac{dn}{100}=0.26dn$. Hence, $G[X]$ contains at least dn/2dn/80.26dn>dn/10dn/2-dn/8-0.26dn>dn/10 edges. Let $X_{0}$ be any superset of $X$ of size $n/C$. Then, $X_{0}$ has average degree at least $Cd/5$, so $G[X_{0}]$ satisfies (i). ∎
As mentioned before, we will repeatedly apply Lemmas˜7.1 and 7.2 to pass to denser neighbourhoods and eventually obtain a large subgraph with constant density. This is our main result for the smallest eigenvalue in this section. However, to make the constants work nicely, we adopt a slightly different proof by considering a subgraph that maximises a carefully chosen potential function. If its density is $o(1)$, Lemmas˜7.1 and 7.2 guarantee a subgraph whose potential function is even higher, which is impossible.
@@
Let $\gamma,\varepsilon,\rho>0$ satisfy $\rho<\frac{1}{2}$, $\varepsilon+6\gamma<1$ and $\frac{\rho}{\varepsilon}+\frac{2\gamma}{1-\varepsilon-4\gamma}<1$. Then there exist $c_{1}=c_{1}(\gamma,\varepsilon,\rho)>0$ and $n_{0}=n_{0}(\gamma,\varepsilon,\rho)$ such that the following holds for every $n>n_{0}$. Let $G$ be an $n$-vertex graph with edge density $p$ and smallest eigenvalue $\lambda_{n}$ such that $p>n^{-\rho}$ and $|\lambda_{n}|\leq n^{\gamma}$. Then $G$ has a subgraph on at least $n^{1-\varepsilon}$ vertices with edge density at least $c_{1}$.